# LAP, Using Action Feasibility for Improved Uncertainty Alignment of Large Language Model Planners

### v2

@@ IV Experiments and Results
We evaluate LAP in a diverse set of language-instructed tasks and environments as described below. Each of these environments is outlined in [3]. We chose to use these environments for easy comparison to [3] and to continue to move towards a standard set of evaluations for the uncertainty alignment task. Each environment has a parameterized scenario distribution with a pre-defined set of possible ambiguities in the human’s instruction. We utilize the truth labels that they provide. The full outline of the dataset distribution and every possible ambiguity is included in the supplementary material. We utilize GPT-4 [36] as the LLM in all experiments and ablate against it separately.
Baselines. Our primary baseline is KnowNo [3], which uses the probability of each option directly from the LLM, and builds a prediction set of options with a probability greater than a threshold value. KnowNo uses conformal prediction to set the threshold value in such a way that a statistical guarantee of a certain success rate is possible. They do nothing to improve the confidence measure itself outside of the task framing as MCQA. All of the additional baselines use this MCQA framing. Other baselines include:
@@
Evaluation Metrics. We primarily use human-help rate and prediction set size versus success rate to compare LAP to baselines. Intuitively, when given a set success rate, these metrics provide information about the help rate and average prediction set size for that success rate, with lower help rates and lower average prediction set sizes better. For a given data sample, the human-help rate is 1 if the prediction set size is greater than 1. Prompt and Binary produce a single success rate and thus a singular help rate and average prediction set size. With KnowNo and LAP, we have a tunable threshold value that is used to confine the prediction set. For these methods, we apply threshold values, $t$, ranging from 0.0000001 to 0.7 for full coverage. The success rate and human help rate or average prediction set size for each threshold value are then plotted, and the area under the resultant curve (AuC) becomes a valuable metric to quantify an improved relationship between task success rate and human help rate.
Figure 4: Comparison of task success rate vs human help rate (left) and prediction set size (right) in our real-world experiments on a ClearPath TurtleBot. Note the improvement over KnowNo across all success rates. LAP with the prompt and context-based A-Feasibility metrics outperforms using each score separately, while all improve upon KnowNo.
@@ IV-A Simulation: Tabletop Rearrangement
In this task, a robot arm is asked to rearrange objects on a table in the PyBullet simulator. Each scenario is initialized with three bowls and three blocks, with one each of blue, green, and yellow, respectively. The task requires the robot to move a certain number of blocks or bowls toward a specific location relative to a different object. For example, “Move the Green Block to the left of the Blue Bowl.” As introduced KnowNow, three different ambiguities are injected into the user instruction: (1) attribute, where ‘block’ or ‘bowl’ is replaced with a potentially ambiguous term like ‘thing’ or ‘cube,’ (2) numeric, where the number of blocks to move is underspecified by using a term like ‘some’ in place of ‘two,’ and (3) spatial, where the exact location like ‘left’ or ‘ front’ is replaced with an ambiguous term like ‘near.’
We show in Figure 3 that LAP with the context-based A-Feasibility metric ($a_{c}$) outperforms KnowNo, as well as all other baselines, with a higher success-to-help ratio across threshold values. Quantitatively, LAP+$a_{c}$ reduces the amount of help needed by over 70%, from 30.1% with KnowNo to 8.9% at a success rate of 60%. Similarly, prediction sets created by LAP are smaller than those from KnowNo. On the Human Help Rate curve, we find that LAP+$a_{c}$ has an Auc of $0.780$ compared to KnowNo with an AuC of $0.699$.
Of Prompt and Binary, both produce a singular success rate and help rate. Binary performed the best for this environment, getting close to the performance of LAP+$a_{c}$ at its success level of 77%. Prompt performs relatively poorly, with a success rate below 60%, illustrating the challenges present when attempting to retrieve the prediction sets directly from the LLM.
LAP with the perception-based A-Feasibility metric, LAP+$a_{perc}$ also under-performs LAP with the context-based score. We believe this is caused by MCQA options, $y^{i}$, becoming more difficult to separate relative to LAP+$a_{c}$, which introduces certainty. We believe that $a_{perc}$ may still be valuable in deployed scenarios without perfect perception. Similarly, LAP+$a_{pr}$ exhibited marginally lower performance than the context-based A-Feasibility metric. We believe this is also due to the context-based scores certainty helping its performance in this simple task.
Figure 5: Example of our real-world experimentation setup. The ClearPath TurtleBot must navigate towards the correct object, at which point a human moves the object into the basket before the robot continues towards a goal location. We chose this setup as no viable mobile manipulator was available to us.
@@ IV-B Real World: Mobile Manipulator in a Kitchen
For this environment, we continue to use the task specifications laid out in KnowNo [3]. Shown in Figure 5, each scenario involves a mobile robot in front of a countertop in an office kitchen, next to a set of recycling/compost/landfill bins. The tasks include picking up objects from the countertop and possibly putting it into a bin, or somewhere else on the countertop. New ambiguities are added relative to the simulator environment, including some involving unsafe actions (e.g. task: “place the bowl in the microwave.”)The full outline of every possible ambiguity is included in the supplementary material. In our lab, no viable mobile manipulator exists, so we run our testing on a ClearPath TurtleBot. During the testing, we manually pick up the object selected by the agent and put it in a basket on the robot, simulating the use of the manipulator. We mark a task a success if the robot indicates the correct actions, moves to within 1 meter of the target object or destination, and orients itself towards the target.
@@
Similar to the simulation environment, we find that LAP outperforms KnowNo and our other baselines. We show these results in Figure 4. Our prompt-based metric and context-based metric combined into a superscore, LAP+$a_{pr}\times a_{c}$ produced our best results, with a human help rate of only 42% relative to KnowNo’s 63% at a 70% planning success rate. These findings are corroborated with LAP having a lower average prediction set size than KnowNo, and an improved AuC of $0.72$ relative to KnowNo’s $0.67$. Qualitatively, we notice that LAP+$a_{pr}\times a_{c}$ improves performance on scenarios with a safety component, likely due to $a_{pr}$ scoring each option partially on its safety. We believe the performance of LAP+$a_{pr}\times a_{c}$ is due to each term contributing differently, with $a_{pr}$ measuring not just possibility, but also safety, for a given action while $a_{c}$ removing many hallucinated objects or destinations from the list of available options entirely. We can identify a similar trend for the Binary and Prompt baselines as in the simulation environmentwith the added complexity and ambiguities of the environment further hampering the Prompt baseline’s ability to produce an accurate prediction set directly.
Different from KnowNo, we ran additional experiments without perfect perception where we found the scene context through the VLM LLaVA. We then produced the perception-based A-Feasibility metric LAP+$a_{perc}$ using the objects listed in the LLaVA output. We only tested this on 20 scenarios due to the manual labor needed to verify LLaVA contexts and subsequent selected options versus the constrained KnowNo scenarios. We found that LAP+$a_{perc}$ was very useful to screen objects that were hallucinated by LLaVA which occurred in 4 of our scenarios. Otherwise, the context-based metric continued to perform better, with all advantaged by the use of the prompt-based metric.
@@
Figure 6: Comparison of LAP+$a_{pr}\times a_{c}$ and KnowNo accross different LLMs. Note that LAP improves upon KnowNo with every model, and that performance generally increases with each model.
@@ IV-C LLM Ablations
To further test LAP, we ablated our model choice, GPT-4 for both our method, LAP+$a_{pr}\times a_{c}$, and KnowNo in our real-world paradigm. Figure 6 shows these results. We show a large improvement over KnowNo across all models. The performance gap between LAP and KnowNo is wider for GPT-3.5-Turbo and GPT-4-Turbo than it is for GPT-4. We hit maximum performance on GPT-4-Turbo at half the human help rate of KnowNo. Additionally, we show that the performance of LAP improves with the model.
As an additional ablation, we also tested LAP using the ablated model for the MCQA generation and scoring phases, but only GPT-3.5-Turbo for finding $a_{pr}$. Performance was largely similar with help rate increasing by under 4% on average. We envision this as a cost-saving measure for users of LAP who want to minimize API or compute costs. These plots can be seen in our supplementary materials.
@@ IV-D Why Not Just Fine-Tune?
In simulation, where we can generate a large set of possible tasks, we attempted to fine-tune gpt-3.5-turbo-1106, the best model available at the time, on each part of LAP. To test generalization, we separate out certain ambiguities of each type, as well as one entire type, for use exclusively in the testing set, which has a distribution identical to our dataset utilized above.
We find that overall help rates increase for the same model by on average 14%. This difference is largely due to impacted performance on ambiguities that were outside of the training data, showing a tendency to overfit seen by others including [3]. Similar issues occurred when fine-tuning on the option scoring component of LAP, decreasing help rates by less than 5% on average, but requiring help more frequently on the withheld ambiguities. This method still underperforms LAP on GPT-4 across most success rates.
In the real-world environment, with even more limited data, fine-tuning the MCQA proposal produced an overfitted model that could not generalize to the testing set of tasks.When finetuning on scoring, help rates decreased on average but with significantly worse performance in withheld categories. The fine-tuned models still underperform LAP on GPT-4. We can draw two separate conclusions from these experiments: (1) Generating a large and diverse enough dataset of scenarios in a robotic paradigm is exceptionally difficult, and (2) The newest, most powerful, models available outperform a fine-tuned older model for this task.
@@ V Conclusions, Limitations, and Future Work
Robots must complete user-provided instructions accurately, with minimal instruction from the user. In this work we introduce LAP, a method of using an action feasibility score to significantly reduce the impact of LLM hallucinations and minimize the amount of human feedback needed when completing tasks. This includes a reduction in human help over prior art KnowNo [3] of 33% at a success rate of 70% in our real-world experiments.
Limitations and Future Work. The primary limitation of the methods we propose is perception. While providing ground-truth perception follows prior art, the uncertainty in perception found in a real-world environment is important to properly handle. An additional limitation is the limited diversity of the dataset we tested on, mainly that it only includes pick-and-place tasks. Future work on more diverse datasets would be valuable.

### v3

@@ IV Experiments and Results
We evaluate LBAP in a diverse set of language-instructed tasks and environments as described below. Each of these environments is outlined in [3]. We chose to use these environments for easy comparison to [3] and to continue to move towards a standard set of evaluations for the uncertainty alignment task. Each environment has a parameterized scenario distribution with a pre-defined set of possible ambiguities in the human’s instruction. We utilize the truth labels that they provide. The full outline of the dataset distribution and every possible ambiguity is included in the supplementary material. We utilize GPT-4 [37] as the LLM in all experiments and ablate against it separately.
Baselines. Our primary baseline is KnowNo [3], which uses the probability of each option directly from the LLM, and builds a prediction set of options with a probability greater than a threshold value. KnowNo uses conformal prediction to set the threshold value in such a way that a statistical guarantee of a certain success rate is possible. They do nothing to improve the confidence measure itself outside of the task framing as MCQA. No other baseline is known to exist which enables asking the user for help in the MCQA paradigm. All of the additional baselines also use this MCQA framing. Other baselines include:
@@
Ablations. To test the necessity of each term in our system, we ablate against the use of the scene grounding, $S$, and world knowledge, $W$, in LBAP. Specifically, the ablations include:
  • LBAP w/ $S$: We evaluate $P(y_{i}|S)$ instead of the full posterior probability. This essentially removes the $P(W|y_{i})$ term.
  • LBAP w/ $W$: We evaluate $P(y_{i}|W)$ instead of the full posterior probability. This essentially removes the $P(S|y_{i})$ term.
The LBAP w/ $S$ ablation is similar to SayCan [1], but using active perception instead of pre-trained value functions for each possible action. The results for these ablations are described in the main experimentation sections. We also ablate against the specific LLM used, changing the LLM for only the MCQA, and fine-tuning the LLM and describe these experiments in a different subsection.
Evaluation Metrics. We primarily use human-help rate and prediction set size versus success rate to compare LBAP to baselines. Intuitively, when given a set success rate, these metrics provide information about the help rate and average prediction set size for that success rate, with lower help rates and lower average prediction set sizes better. For a given data sample, the human-help rate is 1 if the prediction set size is greater than 1. Prompt and Binary produce a single success rate and thus a singular help rate and average prediction set size. With KnowNo and LBAP, we have a tunable threshold value that is used to confine the prediction set. For these methods, we apply threshold values, $t$, ranging from 0.0000001 to 0.7 for full coverage. The success rate and human help rate or average prediction set size for each threshold value are then plotted, and the area under the resultant curve (AuC) becomes a valuable metric to quantify an improved relationship between task success rate and human help rate.
Figure 4: Comparison of task success rate vs human help rate (left) and prediction set size (right) in our real-world experiments on a ClearPath TurtleBot. Note the improvement over KnowNo across all success rates. LBAP improves upon KnowNo and its ablations with just $W$ or $S$ in the Bayesian inference.
@@ IV-A Simulation: Tabletop Rearrangement
In this task, a robot arm is asked to rearrange objects on a table in the PyBullet simulator. Each scenario is initialized with three bowls and three blocks, with one each of blue, green, and yellow, respectively. The task requires the robot to move a certain number of blocks or bowls toward a specific location relative to a different object. For example, “Move the Green Block to the left of the Blue Bowl.” As introduced KnowNow, three different ambiguities are injected into the user instruction: (1) attribute, where ‘block’ or ‘bowl’ is replaced with a potentially ambiguous term like ‘thing’ or ‘cube,’ (2) numeric, where the number of blocks to move is underspecified by using a term like ‘some’ in place of ‘two,’ and (3) spatial, where the exact location like ‘left’ or ‘ front’ is replaced with an ambiguous term like ‘near.’ The scenario distribution is uniform over all possible ambiguities and a testing set is generated for use. All LBAP methods ground the actions using our perception based scoring.
We show in Figure 3 that LBAP outperforms KnowNo, as well as all other baselines and ablations, with a higher success-to-help ratio across threshold values. Quantitatively, LBAP reduces the amount of help needed by over 70%, from 30.1% with KnowNo to 8.9% at a success rate of 60%. Similarly, prediction sets created by LBAP are smaller than those from KnowNo. On the Task Success Rate $\times$ Human Help Rate curve, we find that LBAP has an Auc of $0.780$ compared to KnowNo with an AuC of $0.699$.
Of Prompt and Binary, both produce a singular success rate and help rate. Binary performed the best for this environment, getting close to the performance of LBAP at its success level of 77%. Prompt performs relatively poorly, with a success rate below 60%, illustrating the challenges present when attempting to retrieve the prediction sets directly from the LLM.
LBAP with scene grounding, LBAP w/ $S$ overperformed LBAP without scene knowledge LBAP w/ $W$. We believe this is caused by MCQA options, $y^{i}$, becoming more difficult to separate relative to LBAP w/ $S$, which uses more aggressive probabilities for filtering out hallucinations in from the action options. However, the combined method LBAP with the full Bayesian inference shows the best results, indicating that combining world knowledge and scene information are important for optimal performance.
Figure 5: Example of our real-world experimentation setup. The ClearPath TurtleBot must navigate towards the correct object, at which point a human moves the object into the basket before the robot continues towards a goal location. In this specific example, the perception has grounded a ‘Red Bull,’ a ‘soda,’ and an ‘apple.’ We notice frequent instances of a hallucination caused by the user, for example, the user asking for a ‘fruit’ causing the agent to hallucinate a ‘fruit’ object which is not properly grounded in the scene, instead of the agent selecting the ‘apple.’
@@ IV-B Real World: Mobile Manipulator in a Kitchen
For this environment, we continue to use the task specifications laid out in KnowNo [3]. Shown in Figure 5, each scenario involves a mobile robot in front of a countertop in an office kitchen, next to a set of recycling/compost/landfill bins. The tasks include picking up objects from the countertop and possibly putting it into a bin, or somewhere else on the countertop. New ambiguities are added relative to the simulator environment, including some involving unsafe actions e.g. task: “place the bowl in the microwave” context: In the scene there is a plastic bowl and a metal bowl. The full outline of every possible ambiguity is included in the supplementary material. In our lab, no viable mobile manipulator exists, so we run our testing on a ClearPath TurtleBot. During the testing, we manually pick up the object selected by the agent and put it in a basket on the robot, simulating the use of the manipulator. We mark a task a success if the robot indicates the correct actions, moves to within 1 meter of the target object or destination, and orients itself towards the target.
@@
Similar to the simulation environment, we find that LBAP outperforms KnowNo and our other baselines. We show these results in Figure 4. LBAP indicates with a human help rate of only 42% relative to KnowNo’s 63% at a 70% planning success rate. These findings are corroborated with LBAP having a lower average prediction set size than KnowNo, and an improved AuC of $0.72$ relative to KnowNo’s $0.67$. Qualitatively, we notice that LBAP improves performance on scenarios with a safety component, likely due to the incorporation of world knowledge $W$ evaluating each option partially on its safety. This is backed up by the LBAP w/ $W$ ablation performing better than LBAP w/ $S$. We can identify a similar trend for the Binary and Prompt baselines as in the simulation environment, with Binary outperforming Prompt, but with neither getting particularly close to LBAP. Prompt performs worse relative to the other methods in the real-world environment, exhibiting how the added complexity and ambiguities of the environment further hampering the Prompt baseline’s ability to produce an accurate prediction set directly.
We test LBAP w/ $W$ and LBAP w/ $S$ finding that each improves performance over KnowNo individually. However, combining them in LBAP allows for the greatest effect. We believe this validates our Bayesian inference with each term contributing different things to the final probability. $W$ measures not just possibility, but also safety. Meanwhile, $S$ does a very good job at removing any hallucinated objects or destinations from the list of available options entirely.
@@
Figure 6: Comparison of LBAP and KnowNo across different LLMs. Note that LBAP improves upon KnowNo with every model, and that performance generally increases with each model. We noticed qualitatively that GPT-4-Turbo exhibited poor performance in the MCQA generation phase relative to GPT-4, frequently asking for clarifications instead of producing a set of candidate options, capping the possible success rate.
@@ IV-C LLM Ablations
To further test LBAP, we ablated our model choice, GPT-4 for both our method, LBAP, and KnowNo in our real-world paradigm. Figure 6 shows these results. We show a large improvement over KnowNo across all models. The performance gap between LBAP and KnowNo is wider for GPT-3.5-Turbo and GPT-4-Turbo than it is for GPT-4. We hit maximum performance on GPT-4-Turbo at half the human help rate of KnowNo. Additionally, we show that the performance of LBAP improves with the model. This encouraging behavior implies that when models get released in the future, LBAP will continue to apply and increase their performance for these tasks.
As an additional ablation, we also tested LBAP using the ablated model for the MCQA generation and scoring phases, but only GPT-3.5-Turbo for evaluating the world knowledge. Performance was largely similar with help rate increasing by under 4% on average. We envision this as a cost-saving measure for users of LBAP who want to minimize API or compute costs. These plots can be seen in our supplementary materials.
@@ IV-D Why Not Just Fine-Tune?
In simulation, where we can generate a large set of possible tasks, we attempted to fine-tune gpt-3.5-turbo-1106, the best model available at the time, on each part of LBAP. To test generalization, we separate out certain ambiguities of each type, as well as one entire type, for use exclusively in the testing set, which has a distribution identical to our dataset utilized above.
We find that overall help rates increase for the same model by on average 14%. This difference is largely due to impact on performance for ambiguities that were outside of the training data, showing overfitting as seen by others including [3]. Similar issues occurred when fine-tuning on the option scoring component of LBAP, decreasing help rates by less than 5% on average, but requiring help more frequently on the withheld classes. This method underperforms LBAP on GPT-4 across most success rates. However, prediction set sizes were decreased more than LBAP on average, showing the model improving its rejection of poor options.
In the real-world environment, with even more limited data, fine-tuning the MCQA proposal produced an overfitted model that could not generalize to the testing set of tasks.When fine-tuning on the scoring task, we saw a similar phenomenon to in the simulation paradigm with improved performance on samples inside the training distribution, but poorer performance on those outside. When finetuning on scoring, help rates decreased on average but with significantly worse performance in withheld categories. The fine-tuned models still underperform LBAP on GPT-4. We can draw two separate conclusions from these experiments: (1) Generating a large and diverse enough dataset of scenarios in a robotic paradigm is exceptionally difficult, and (2) The newest, most powerful, models available outperform a fine-tuned older model for this task.
@@ V Conclusions, Limitations, and Future Work
Robots must complete user-provided instructions accurately, with minimal instruction from the user. In this work we introduce LBAP, a method of using Bayesian inference to significantly reduce the impact of LLM hallucinations and minimize the amount of human feedback needed when completing tasks. This includes a reduction in human help over prior art KnowNo [3] of 33% at a success rate of 70% in our real-world experiments. LBAP presents a strong improvement over the confidence score used in KnowNo in all cases while retaining the ability to use conformal prediction to set up statistical performance guarantees.
Limitations and Future Work. The primary limitation is the limited diversity of the dataset we tested on, mainly that it only includes pick-and-place tasks. Future work on more diverse datasets would be valuable. This would especially help to validate the incorporation of world knowledge to tasks that involve rules or restrictions that are not otherwise obvious but could be incorporated into the prompts.